# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@
Abstract.
Malware detection is one of the most ubiquitous applications of machine learning (ML) in security. Prior ML-based detectors can consume program features extracted statically from binary code or dynamically from execution traces. Dynamic detectors most commonly rely on executing programs in synthetic environments (sandboxes) that allow controlled malware analysis. However, sandbox traces are unavailable to the last line of defense offered by security vendors: real-time malware detection at endpoints. A detector at endpoints consumes the traces of programs running on hosts (i.e., real-world environments), as sandbox analysis, typically done in the cloud, might introduce intolerable delays. Despite their success in the sandbox setting, research hints at potential challenges for ML methods at endpoints, e.g., highly variable malware behaviors in the wild. Nonetheless, the impact of these challenges on existing approaches and how their excellent sandbox performance translates to the endpoint scenario remain unquantified.
We present the first measurement study of the performance of ML-based malware detectors at real-world endpoints. Leveraging a dataset of sandbox traces we collected and a dataset of in-the-wild program traces from Avllazagaj et al. (avllazagaj2021malware, ); we evaluate two scenarios where the endpoint detector was trained on (i) sandbox traces (convenient and accessible); and (ii) endpoint traces (less accessible due to needing to collect telemetry data). This allows us to identify a wide gap between prior methods’ sandbox-based detection performance—over 90%—and endpoint performance—about 20% (scenario (i)) to 50% (scenario (ii)). We pinpoint and characterize the challenges contributing to this gap, including label noise, behavior variability, sandbox evasion, or environment-specific features. Building on this, we suggest strategies to close the gap, yielding a relative improvement of 5–30% over the baselines. Our evidence suggests that applying detectors trained on sandbox data to endpoint detection is challenging. The most promising direction is training detectors directly on endpoint data, which marks a departure from current practice. To promote research, we implement a leaderboard for realistic detector evaluations—a format that has fueled progress in many ML tasks.
malware detection, dynamic analysis, machine learning
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Security and privacy Malware and its mitigation
1. Introduction
Detection of malware threats is crucial for governments, enterprises, and end users as there are significant (and growing) financial and safety harms of malware infections (reddick_2022, ), which has created a \$7 Billion industry in 2022 with many players (imarcreport, ). Malware detection appears to be remarkably effective: industry-standard evaluations of commercial anti-malware products (avcomparisons, ), and prior research on malware detection using machine learning (ML) methods (raff2018malware, ; karbab2019maldy, ), routinely report that over 99% of malware samples in a standard corpus can be detected, with very few false-positives.
@@ 1. Introduction
Sandbox traces Sandbox traces Standard classifier $\sim$95(karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, )
@@
Table 1. Performance of ML-based behavioral malware detection under different settings.
@@
Malware detection is often performed at endpoints where an endpoint security solution (quarantine, ) monitors host devices to detect threats in real time. In practice, these solutions rely on a chain of techniques (vmray, ), including static analysis and dynamic analysis. Conventional static methods, such as blocklists, reputation systems, and signatures, operate without executing the program. As static methods can be bypassed through obfuscation or polymorphism (moser2007limits, ; song2007infeasibility, ), dynamic analysis has become standard, offered by most vendors (kasperskyrealtime, ).
@@
Dynamic analysis relies on observing the execution behaviors of a sample to detect whether it displays malicious behaviors. The typical paradigm for obtaining such behavior-based detectors is to detonate (execute) a large set of known samples (malware and benign) in a controlled environment (a sandbox), collect their execution traces, and learn to separate malicious and benign behaviors. Several ML models have demonstrated exceptional performance in this task (mariconti2016mamadroid, ; karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ). Unfortunately, sandbox traces are unavailable to models in endpoint detection as detonation cannot be done in real time, despite advances such as on-premise sandboxes (onpremsandbox, ). Consequently, these models must rely on traces observed on real-world hosts to detect malware.
@@
Research has outlined several challenges associated with this task. First, program behaviors are environment-sensitive, and, especially for malware, two traces of the same sample from two different hosts often have little overlap (avllazagaj2021malware, ). As a result, a model trained on a trace captured on one host might fail to generalize to other hosts. Second, techniques frequently employed to evade sandbox analysis (lindorfer2011detecting, ; balzarotti2010efficient, ; galloro2022systematical, ) can mislead models trained on sandbox traces into learning features that do not capture actual malware behaviors. Although these challenges were acknowledged in prior work (barecloud, ; kirat2015malgene, ; jindal2019neurlux, ; avllazagaj2021malware, ), we lack understanding of how they manifest, their implications for endpoint detectors, and potential solutions.
@@
Our work conducts the first quantitative study into the efficacy of ML-based methods for endpoint detection to demystify these challenges. We specifically focus on two scenarios: (i) an endpoint detector trained using only sandbox traces and (ii) an endpoint detector trained using real-world endpoint traces. Existing techniques can conveniently gather large-scale datasets of sandbox traces, making (i) a practical and accessible option. On the other hand, collecting endpoint traces is attainable mainly by vendors who receive telemetry data from the wild as it is still an unsolved problem to simulate such data in controlled environments (such as sandboxes) at scale (schreuders2017security, ; miramirkhani2017spotless, ; mills2020investigating, ). These requirements make (ii) a less accessible but likely a more successful option. Consequently, both scenarios are relevant in practice, which motivates us to study their challenges.
@@
We employ three deep-learning-based approaches (karbab2019maldy, ; jindal2019neurlux, ; trizna2023nebula, ) that can achieve $\sim$95true-positive rate (TPR) at 1% false-positive rate (FPR) when evaluated on sandbox traces. We use a dataset from Avllazagaj et al. (avllazagaj2021malware, ) of around 1M endpoint traces from over 25K samples, recorded on real-world hosts by a commercial anti-malware vendor. The malicious samples in this dataset were undetected by the vendor’s defenses at that time, corresponding to real-world infections. Consequently, this data reflects the realistic threats that behavior-based detectors must combat. Additionally, to study scenario (i), we collect a dataset of traces (contemporaneous to our endpoint dataset) from two sandboxes.
@@
Our initial finding is that ML methods exhibit lower performance in both scenarios (i) ($\sim$17TPR) and (ii) ($\sim$49TPR), in contrast with their excellent performance on sandbox traces (see Table 1). We study the low performance in (i) from the lens of distribution shift, a problem that plagues ML in many applications (koh2021wilds, ). Others in the security community have articulated the challenge of concept drift, where the data distribution shifts over time (jordaney2017transcend, ). We explore whether the problem can be addressed using existing tools for concept drift. However, we found that existing concept drift methods are not enough to address the differences in data distribution between sandbox vs endpoint traces: we discovered that endpoint traces of benign samples are detected as “drifting” (whereas in prior work concept drift has mostly affected malware (barbero2022transcending, )), and rejecting drifting samples does not improve the performance enough.
@@
Investigating the root causes of the low performance, we discovered that endpoint detectors are applied to a different distribution of samples than considered in prior work: they are typically applied only to the hardest-to-classify samples. In past research, researchers have trained and evaluated detectors on a corpus of samples from repositories (karbab2019maldy, ; jindal2019neurlux, ; dambra2023decoding, ) such as MalwareBazaar (bazaar, ). However, endpoint malware detection systems employ a pipeline, where samples are first categorized using basic methods (e.g., signatures, static analysis), and ML-based classifiers are only applied to samples that cannot be resolved using basic methods (vmray, ). Consequently, in the wild, endpoint classifiers are only applied to samples that tend to be harder to classify than those found in a standard corpus. Prior research has not considered this factor, and we find that it causes a significant drop in performance: it reduces the performance of classifiers trained on sandbox traces from 95% (for samples from a standard corpus) to $\sim$60TPR (when the distribution of samples is adjusted to take into account earlier stages in the pipeline). This shows that prior evaluations of behavior-based detectors have greatly over-estimated their effectiveness in the wild.
@@
Second, we study the impact of variable program behaviors across different environments. We measure large differences between a sample’s sandbox trace and its endpoint traces. Sandbox traces lack diversity: collecting multiple traces by running a group of related samples, e.g., from the same malware family, in a sandbox yields very similar traces. This introduces spurious features that do not generalize to other environments. Conversely, endpoint traces are highly diverse, making a model’s predictions on different traces of the same sample inconsistent, which hurts detection performance. Moreover, we provide the first evidence that sandbox-evading malware (40–80% of all malware (galloro2022systematical, )) skews a model in scenario (i) towards classifying very short traces (often an indication of evasion (kuchler2021does, )) as malware. This correlation is spurious as it is absent in endpoint traces. Thus, this correlation causes prior evaluations using sandbox traces to over-estimate the accuracy of endpoint classifiers.
@@
Our characterization of these challenges allows us to pinpoint avenues for improvement in endpoint detection. To improve the performance on the distribution of difficult-to-classify samples, we explore soft-labeling (effective against label noise (lukasik2020does, )) and using more accurate distributions for training. Against variable behaviors, we employ a technique popular in other areas of ML: invariant learning (zhao2019learning, ). In particular, we train our models to make consistent predictions on different traces of the same sample to prevent them from learning unreliable, environment-specific features. When combined, these strategies lead to moderate but promising gains, from 17% to 22% TPR in scenario (i) (30% relative improvement) and from 49% to 52% TPR in scenario (2) (5% relative improvement).
@@
We believe our results should serve as a call to action for the research community. Previous research has suggested that it is possible to achieve 95% TPR, which might leave an impression that progress is saturated, there is not much room for further improvement, and the problem of ML-based malware detection is more or less solved. Our results show that the reality is different: the problem is not solved, the actual performance on malware in the wild is far worse, and there are major unsolved challenges and significant room for further improvement. To stimulate further research on this challenging problem, we have designed a leaderboard (malwaredetectioninthewild.github.io (leaderboard, )), a popular format in ML (croce2021robustbench, ), for realistic evaluations of behavior-based malware detectors. Participants receive access to our sandbox dataset and all our metadata and can submit their detectors, which we will rank according to their endpoint performance. This will hopefully allow the community to identify promising directions.
Contributions. (I) We measure the performance discrepancy of ML-based approaches between sandbox-only and endpoint settings (Section 5). (II) We identify the challenges in endpoint detection to demystify this discrepancy and propose improvements (Section 6 and Section 7). (III) We organize a leaderboard for realistic evaluations of detectors (Section 8).
2. Background and Related Work
@@
Dynamic Malware Analysis. Most work in dynamic analysis focuses on analyzing the behaviors of a sample detonated in controlled environments, such as sandboxes (willems2007toward, ; bayer2009view, ). As dynamic analysis has become a staple, malware has started including checks that suppress malicious activities if the environment is fingerprinted as a sandbox, known as evasive malware (lindorfer2011detecting, ). Although many strategies have been developed to prevent fingerprinting (liu2022enhancing, ), this is still an ongoing arms race (miramirkhani2017spotless, ). Although researchers have explored methods to analyze samples in bare metal environments (balzarotti2010efficient, ; barecloud, ), they fall short in preventing evasion (miramirkhani2017spotless, ). To our knowledge, we have taken the first step toward measuring the implications of evasion for ML-based detectors at endpoints. A recent large-scale measurement study over variable program behaviors in the wild has found that a given malware sample can behave significantly differently across time and in different real-world hosts (avllazagaj2021malware, ). Our work connects to this line of work as we are interested in quantifying the impact of these challenges, such as evasion or variability, on a malware detector deployed for endpoint detection at hosts in the wild.
@@
ML for Malware Detection. ML-based methods are widely used in research and practice with static (raff2018malware, ) and dynamic features (huang2016mtnet, ). In detection with dynamic features, methods that essentially treat a program’s execution trace as a natural language document (a sequence of tokens) and adapt off-the-shelf methods from the ML community have been successful (miller2016reviewer, ; karbab2019maldy, ; jindal2019neurlux, ; kuchler2021does, ; trizna2023nebula, ). Most of these methods are trained and evaluated on the traces from a pre-configured sandbox on samples collected from public repositories; some popular ones are listed here (awesomemalware, ). We aim to understand the implications of these practices and the efficacy of popular ML approaches for detecting malware in real-world hosts with dynamic features.
@@
Distribution Shifts in ML. Distribution shift occurs when a model’s training and testing distributions have significant differences, which hurts the performance (koh2021wilds, ). For example, as new malware variants emerge and old ones disappear over time, the performance of a detector that has not been kept up-to-date will deteriorate, known as concept drift (pendlebury2019tesseract, ). A popular paradigm to deal with concept drift is deploying drift detectors to reject samples that would have been misclassified and training on them later to update the model (jordaney2017transcend, ; barbero2022transcending, ; chen2023continuous, ). We deploy a drift detector to characterize the differences between the distribution shifts in our problem and concept drift. Although there is no silver bullet, the ML community has also proposed many ideas to tackle distribution shifts in different contexts, such as learning domain invariant features (ganin2015unsupervised, ; 2013gongconnecting, ), distributionally robust optimization (sagawa2020investigation, ), or continuous learning (chen2023continuous, ). We adapt some of these ideas, such as invariant learning, to our problem to measure their benefits for malware detection at endpoints.
@@
Limitations of ML for Security. Research suggests that popular ML methods have many pitfalls when applied to security tasks (arp2020and, ). For example, in malware detection, most work has been found to overestimate the success of the methods due to impossible time splits of training and testing sets (pendlebury2019tesseract, ). A line of work also focuses on how ML methods successful in lab-only evaluations break down when they are deployed in the real world due to distribution shifts, for example, in the context of network anomaly detection (jacobs2022ai, ), website fingerprinting (cherubin2022online, ) or malware detection (das2019sok, ; aghakhani2020malware, ). In these contexts, models are known to learn spurious features, such as specific IP addresses (jacobs2022ai, ), that are artifacts of the experimental setup in the lab and do not apply to realistic settings. Our work also investigates the limitations of ML methods in malware detection with dynamic features when they are trained and evaluated in controlled settings (e.g., using sandboxes) and deployed in the wild.
3. Endpoint Malware Detection

### v2

# ML-Based Behavioral Malware Detection Is Far From a Solved Problem

@@
Abstract
Malware detection is a ubiquitous application of Machine Learning (ML) in security. In behavioral malware analysis, the detector relies on features extracted from program execution traces. The research literature has focused on detectors trained with features collected from sandbox environments and evaluated on samples also analyzed in a sandbox. However, in deployment, a malware detector at endpoint hosts often must rely on traces captured from endpoint hosts, not from a sandbox. Thus, there is a gap between the literature and real-world needs.
We present the first measurement study of the performance of ML-based malware detectors at real-world endpoints. Leveraging a dataset of sandbox traces and a dataset of in-the-wild program traces, we evaluate two scenarios: (i) an endpoint detector trained on sandbox traces (convenient and easy to train), and (ii) an endpoint detector trained on endpoint traces (more challenging to train, since we need to collect telemetry data). We discover a wide gap between the performance as measured using prior evaluation methods in the literature—over 90%—vs. expected performance in endpoint detection—about 20% (scenario (i)) to 50% (scenario (ii)). We characterize the ML challenges that arise in this domain and contribute to this gap, including label noise, distribution shift, and spurious features. Moreover, we show several techniques that achieve 5–30% relative performance improvements over the baselines. Our evidence suggests that applying detectors trained on sandbox data to endpoint detection is challenging. The most promising direction is training detectors directly on endpoint data, which marks a departure from current practice. To promote progress, we will facilitate researchers to perform realistic detector evaluations against our real-world dataset.
I Introduction
Detection of malware threats is crucial for governments, enterprises, and end users as there are significant (and growing) financial and safety harms of malware infections [1], which has created a \$7 Billion industry in 2022 with many players [2]. Malware detection appears to be remarkably effective: industry-standard evaluations of commercial anti-malware products [3], and prior research on malware detection using machine learning (ML) methods [4, 5], routinely report that over 99% of malware samples in a standard corpus can be detected, with very few false positives.
  This work has been accepted for publication in the IEEE Conference on Secure and Trustworthy Machine Learning (SaTML). The final version will be available on IEEE Xplore.
@@ I Introduction
Sandbox traces Sandbox traces Standard classifier $\sim$95[5, 6, 7]
@@
TABLE I: Performance of ML-based behavioral malware detection under different settings.
@@
Malware detection is often performed at endpoints, where an endpoint security solution [8] monitors host devices to detect threats. In practice, these solutions employ a chain of techniques [9], including static analysis and dynamic analysis. Static methods, such as blocklists and signatures, operate without executing the program. As static methods can readily be bypassed via obfuscation or polymorphism [10, 11], dynamic analysis has become a standard offering [12].
@@
Dynamic analysis relies on observing the execution behaviors of a sample to detect whether it displays malicious behaviors. The typical paradigm for obtaining a behavioral detector is to detonate (execute) a large set of known samples (malware and benign) in a controlled environment (a sandbox), collect their execution traces, and learn to separate malicious and benign behaviors. Several ML models have been effective in this task [13, 5, 6, 7]. Unfortunately, endpoint detection models cannot use sandbox traces to decide whether a sample under observation is malicious, as detonation cannot be done in real time, despite advances such as on-premise sandboxes [14]. Consequently, these models must rely on traces observed at real-world hosts to detect malware.
@@
Research has outlined several challenges associated with this task. First, program behaviors are environment-sensitive, and, especially for malware, two traces of the same sample from two different hosts often have little overlap [15]. As a result, an ML model trained on a trace captured on one host might fail to generalize to others. Second, techniques frequently employed to evade sandbox analysis [16, 17] can mislead models trained on sandbox traces into learning spurious features that do not capture actual malware behaviors. Although these challenges have been previously articulated in the research literature [18, 19, 6, 15], we lack an understanding of how they manifest and can be addressed.
@@
Our work conducts the first quantitative study into the efficacy of ML-based methods for endpoint detection to demystify these challenges. We specifically focus on two scenarios: (i) an endpoint detector trained using only sandbox traces and (ii) an endpoint detector trained using real-world endpoint traces. Existing techniques can conveniently gather large datasets of sandbox traces, making (i) a practical and accessible option. On the other hand, collecting endpoint traces is attainable mainly by vendors who receive telemetry data from the wild, as simulating such data at scale in controlled environments (such as sandboxes) is still an open problem [20, 21, 22]. This makes (ii) a less accessible but likely a more effective option. Consequently, both scenarios are relevant in practice, which motivates us to study their challenges.
@@
We evaluate three ML approaches [5, 6, 7] that reach 95% true-positive rate (TPR) at 1% false-positive rate (FPR) when evaluated on sandbox traces. We use a dataset from Avllazagaj et al. [15] containing around 1M endpoint traces from 26K samples, recorded on real-world hosts by an anti-malware vendor. The malicious samples in this dataset were undetected by the vendor’s defenses at that time and caused real-world infections. Consequently, this data reflects the realistic threats that behavioral detectors must combat. Additionally, we collect a dataset of traces (contemporaneous to our endpoint data) from two sandboxes to realistically study scenario (i).
@@
Our initial finding is that ML methods perform poorly in both scenarios (i) ($\sim$17TPR) and (ii) ($\sim$49TPR), in contrast with their excellent performance on sandbox traces (see Table I). We study the low performance in (i) from the lens of distribution shift, a problem that plagues ML in many applications [23]. Similarly, the security community has articulated the challenge of concept drift, where the data distribution shifts over time [24]. We attempt to address the differences in data distribution between sandbox vs. endpoint traces using classification with rejection (an existing tool against concept drift [24]) and found that it does not improve the performance enough. Interestingly, endpoint traces of benign samples are also rejected as “drifting” in high proportions, whereas in prior work concept drift has mostly affected malware [25].
@@
Investigating why existing methods perform poorly, we first discovered that endpoint detectors are applied to a different distribution of samples than considered in prior work: they are typically applied only to the hardest-to-classify samples. Past research has trained and evaluated detectors on a corpus of samples from repositories [5, 6, 26] such as MalwareBazaar [27]. However, endpoint malware detection systems employ a pipeline where basic methods (e.g., static signatures) attempt to categorize samples first, and ML-based classifiers are applied only to samples that remain unresolved [9]. Consequently, in the wild, endpoint classifiers are only applied to samples that tend to be harder to classify than an average sample in a standard corpus. Prior research has overlooked this factor, and we find that it causes a significant drop in performance: it reduces the performance of classifiers trained on sandbox traces from 95% (for samples from a standard corpus) to $\sim$60TPR (when we adjust the distribution of samples to take into account earlier stages in the pipeline). This shows that prior evaluations have vastly overestimated the effectiveness of behavioral detectors in the wild.
@@
Second, we study the impact of variable program behaviors across different environments. We discovered large differences between a sample’s sandbox trace and its endpoint traces. Sandbox traces lack diversity: collecting multiple traces by running a group of related samples, e.g., from the same malware family, in a sandbox produces very similar traces. This introduces spurious features that do not generalize to other environments. In contrast, endpoint traces are diverse, which hurts performance by making a model’s predictions on different traces of the same sample inconsistent. Moreover, we provide the first evidence that sandbox-evading malware (40–80% of all malware [17]) biases a model in scenario (i) to classify very short traces (often an indication of evasion [28]) as malware. This correlation is spurious as it is absent in endpoint traces; thus, it causes prior evaluations using sandbox traces to overestimate the accuracy of endpoint classifiers.
@@
Characterizing these challenges allows us to explore avenues for performance improvements. To improve the performance on the distribution of difficult-to-classify samples, we use soft labeling (effective against label noise [29]) and more accurate distributions for training. Against variable behaviors, we employ a technique popular in other areas of ML: invariant learning [30]. In particular, we train our models to make consistent predictions on different traces of the same sample, forcing them to learn robust, environment-independent features. Together, these strategies lead to moderate gains, from 17% to 22% TPR in scenario (i) and from 49% to 52% in scenario (ii) (30% and 5% relative improvement, respectively).
@@
We believe our results should serve as a call to action for the community. Prior research has suggested that it is possible to achieve over 95% TPR, which might leave the impression that progress is saturated and there is not much room for improvement in behavioral malware detection using ML. We show that the reality is different: the problem is not solved, the actual performance on malware in the wild is far worse, and there are major unsolved challenges and significant room for further improvement. Moreover, ML methods that have shown success on standard benchmarks, such as group robustness [31], struggle due to the complexities of this domain. We plan to stimulate progress on this problem by releasing our sandbox dataset and metadata and offering a pipeline that allows researchers to evaluate their behavioral malware detectors against our real-world endpoint data. Our public website (https://malwaredetectioninthewild.github.io/) includes the details on this evaluation pipeline and data release.
Contributions. (I) We measure the performance discrepancy of ML-based malware detectors between sandbox-only and endpoint settings. (II) We characterize the ML challenges, such as distribution shifts, behind this discrepancy. (III) We explore ML methods to improve the endpoint performance.
II Background and Related Work
@@
Dynamic Malware Analysis. Most work in dynamic analysis focuses on analyzing the behaviors of a sample detonated in controlled environments, such as sandboxes [32, 33]. As dynamic analysis has become a staple, malware has started including checks that suppress malicious activities if the environment is fingerprinted as a sandbox, known as evasive malware [34]. Although many strategies have been developed to prevent fingerprinting [35], this is still an ongoing arms race [21]. Even the methods that analyze samples in bare metal environments [16, 18] struggle to prevent evasion [21]. To our knowledge, we have taken the first step toward measuring the implications of evasion for ML-based detectors at endpoints. A recent large-scale measurement study over variable program behaviors in the wild has found that a given malware sample can behave significantly differently across time and in different real-world hosts [15]. Our work connects to this line of work as we are interested in quantifying the impact of these challenges, such as evasion or variability, on a malware detector deployed for endpoint detection at hosts in the wild.
@@
ML for Malware Detection. ML-based methods are widely used in research and practice with static [4] and behavioral features [36]. In behavioral detection, methods that treat a program’s execution trace as a text document (a sequence of tokens) and adapt existing ML architectures have shown promise [5, 6, 7]. Most of these methods are trained and tested on the traces from a pre-configured sandbox for samples collected from public repositories (see [37] for the popular ones). We aim to understand the implications of these practices and the efficacy of popular ML approaches for detecting malware with behavioral features in real-world hosts.
@@
Distribution Shifts in ML. Distribution shift occurs when a model’s training and testing distributions have significant differences, which hurts the performance [23]. For example, as new malware variants emerge and old ones disappear over time, the performance of a detector that has not been kept up-to-date will deteriorate, known as concept drift [38]. An effective way to deal with concept drift is deploying drift detectors to reject samples that would have been misclassified and training on them later to update the model [24, 25, 39]. We deploy a drift detector to characterize how the distribution shifts in our problem differ from concept drift. The ML community has also proposed other ideas to tackle distribution shifts in different contexts, such as domain invariant features [40, 41], distributionally robust optimization [31], or continuous learning [39]. We adapt some of these ideas, such as invariant learning, to assess their benefits in our problem.
@@
Limitations of ML for Security. Research suggests that popular ML methods have many pitfalls when applied to security tasks [42]. For example, in malware detection, the success of many methods has been overestimated due to impossible time splits of training and testing sets [38]. Another line of work focuses on how ML methods successful in lab-only evaluations break down in the real world due to distribution shifts, e.g., in the context of network anomaly detection [43], website fingerprinting [44] or malware detection [45, 46]. In these contexts, models are known to learn spurious features, such as specific IP addresses [43], that are artifacts of the experimental setup and do not apply to realistic settings. We investigate the limitations of ML methods specifically in behavioral malware detection when they are trained and evaluated in controlled settings (e.g., using sandboxes) but deployed in the wild.
III Endpoint Malware Detection